OpenStudy (anonymous):

The end points of line segment PQ are P(1, 3), and Q(7, 7). What is the midpoint of line segement PQ?

OpenStudy (hoblos):

x= (1+7)/2 = 4 y=(3+7)/2 = 5 (4,5)